# SADG: Segment Any Dynamic Gaussian Without Object Trackers

### v1

@@ 6.3 Google Immersive
Ours (2%) | 0.8719 | 0.9826 | 0.4068 | 0.8368 | 0.9790 | 2.3117 | 0.8738 | 0.9923 | 2.2506 | 0.9154 | 0.9942 | 9.6083 | 0.9258 | 0.9904 | 1.6118 |
SADG (10%) | 0.8707 | 0.9828 | 1.5782 | 0.8384 | 0.9794 | 11.8430 | 0.8792 | 0.9930 | 6.1878 | 0.9160 | 0.9939 | 71.2429 | 0.9283 | 0.9908 | 6.8211 |
SADG (25%) | 0.8762 | 0.9832 | 3.9425 | 0.8461 | 0.9809 | 43.2682 | 0.8886 | 0.9935 | 18.8976 | 0.9180 | 0.9943 | 221.7642 | 0.9297 | 0.9907 | 19.0767 |
@@ 7 Testing Implementation Details
After the dynamic geometry reconstruction and the Gaussian features are properly trained, we proceed to our GUI to select the object of interest and evaluate the performance of the model by mIoU and mAcc as discussed in Sec. 4.4.
@@
HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | |||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SA4D [16] | 0.7767 | 0.9633 | 0.8832* | 0.9391* | 0.7987* | 0.9835* | 0.8271* | 0.9734* |
SADG (Ours) | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
@@ 8 Ablation Studies
8.1 Filtering
We noticed that the raw segmented object from the selected Gaussian clusters sometimes has some artifacts near the outline of the objects as illustrated in Fig. 9. Therefore, we adopt a simple yet effective filtering step to filter out those artifacts. The procedure of the filtering is as follows:
In each selected Gaussian cluster , we compute the mean Gaussian features in the selected cluster .
We compute the per-Gaussian similarity score between the Gaussian features inside the selected cluster and mean Gaussian features of the selected cluster with cosine similarity.
Delete Gaussians that have lower similarity scores than the threshold .
As shown in Tab. 2, the proposed filtering step can effectively improve the average mIoU and mAcc for each benchmark.
@@ 8.2 Smooth Gaussian Features
In Sec. 3.2 we introduced smooth Gaussian features (Eq. 6). We provide a detailed analysis of the selection of the number of nearest neighbors for the KNN algorithm in Tab. 3. Overall, when smooth Gaussian features are enabled ( and ), the performance is better than when smooth Gaussian features are disabled (). Since there is no significant improvement between and , we eventually chose for our pipeline to avoid additional overheads with the KNN sampling.
@@ 8.3 Gaussian Feature Clustering
In Sec. 3.2 we additionally introduced Gaussian Feature Clustering and the sub-sampling strategy to run the clustering in real-time with scenes that normally contain more than 300k Gaussians. We perform time analysis with different sub-sampling numbers of Gaussians (2%, 10%, and 25%) on RTX3080 and Intel i7-12700. For each scene, we compute the time for clustering using Python timeit package and take the average clustering time for five runs per scene. We report the average clustering time across all scenes in each benchmark and the corresponding mIoU and mAcc for our object of interest in Tab. 4. Overall, with different sub-sampling of Gaussians (2%, 10%, and 25%), the mIoU and mAcc stay almost unchanged. However, as increasing the number Gaussians results in a longer clustering time, we believe that 2% is the optimal sub-sampling number for clustering.
@@
There are various clustering techniques and K-means [29, 32, 15, 39] is one of the most widely used. K-means partitions data into clusters, where each feature belongs to the cluster with the nearest centroid (mean of the cluster). The value must be predefined, and the algorithm iteratively adjusts the centroids to minimize the sum of squared distances between features and their corresponding cluster centers. Therefore, the performance of the algorithm heavily depends on the optimal value of .
@@
8.5 Time and Storage Analysis
We performed a comprehensive time analysis of SADG using an RTX 3080 GPU and an Intel i7-12700 CPU on the NeRF-DS [52] dataset. To evaluate its performance, we compared it with DGD [25], the only publicly available model addressing segmentation in dynamic novel views. The results are summarized in Tab. 6.
Most existing approaches for novel view segmentation [57, 6, 16] require precomputing SAM masks or object masks prior to training. In contrast, DGD dynamically computes DINOv2 feature maps for supervision during training. For a fair comparison, we precomputed the DINOv2 feature maps for all training images in DGD and stored them beforehand on an SSD, alongside our precomputed SAM masks.
As shown in Tab. 6, DGD [25] spends little preprocessing time for DINOv2 features (approximately 5 seconds for all training images per sequence). Although our SADG requires preprocessing time to generate SAM masks, this process only needs to be done once before training. DGD assigns a 384-dimensional semantic feature to each Gaussian. Rendering these high-dimensional features into 2D is computationally inefficient, significantly slowing down the training process. SADG uses 32-dimensional Gaussian features, which can be rendered much more efficiently. As a result, SADG achieves over 3 faster total time (preprocessing + training) than DGD while only requiring 4.5 less space in the storage.
In addition, we evaluated the processing time for click prompts in our GUI. For smaller scenes with less than 200k Gaussians such as as_novel_view, segmenting an object via a click prompt takes approximately 0.05 seconds. For larger scenes with more than 1,000k Gaussians captured by a multi-view camera setup, such as 01_Welder, the processing time per click prompt is around 0.4 seconds.
Avg. Preprocessing Time (min) | Avg. Training Time (min) | Total Time (min) | Avg. Size of Gaussians (MB) | |
DGD [25] | 0.08 | 103 | 103 | 285 |
SADG (Ours) | 16 | 16 | 32 | 60 |
9 Comparison Against Segment Any 4D Gaussians Model
We provide additional quantitative and qualitative comparisons with the concurrent work on 4D segmentation and editing, SA4D [16]. The pipeline consists of three stages: (1) dynamic 3D reconstruction based on 4D-GS model [50], (2) identity feature field learning, and (3) post-processing with outlier removal. Since the authors have not released the official code, we use the unclean version from https://github.com/guanjunwu/sa4d for training and testing the model on our benchmark. We do not provide the performance on NeRf-DS-Mask with specular and transparent moving objects in Tab. 5 since 4D-GS fails to converge on most of the sequences apart from two (basin_novel_view, press_novel_view). Similarly, we exclude two sequences from the HyperNeRF-Mask dataset, such as hand1-dense-v2 and espresso due to 4D reconstruction failure.
In quantitative assessment, we select the top 30 segmentation results to more closely reflect the test set used in the original SA4D benchmark (not released). We also uniformly sub-sample point cloud (100K points) from SfM [45]. As seen in Tab. 5, our SADG significantly outperforms SA4D on all datasets. From the qualitative results in Fig. 12 and Fig. 13, we notice that rendered masks are larger than the underlying object and include a lot of unrelated regions even after the outlier removal in the post-processing step, which negatively affects objects’ segmentation accuracy. We further demonstrate some editing comparisons in Fig. 17, Fig. 16, and Fig. 15. Due to inaccurate segmentation, composition, and removal, SA4D results contain many spurious regions, degrading the editing quality.
10 Quantitative Results per Sequence
For completeness to the average numbers per datasets (Tab. 1), we report the performance of our method individually on all benchmarks in Tab. 7, Tab. 8, Tab. 9, Tab. 10, and Tab. 11.
11 Segmentation Benchmarks
12 Additional Qualitative Results
We provide additional qualitative results in Fig. 12 and Fig. 13 to further showcase the effectiveness of SADG over other methods. While SAGA [7] suffers from degenerate Gaussians, our reconstruction backbone and effective post-processing mitigate the issue. At the same time, rendered masks from other methods, such as DGD [25] and SA4D [16], are inaccurate and introduce spurious regions. Our method demonstrates the best performance among all approaches.
13 Further Scene Editing Examples
We present further scene composite examples of SADG in Fig. 14. We also compare the quality of object removal and scene composition with SA4D [16] in Fig. 16 and Fig. 17. Overall, as discussed in Sec. 9, SADG outperforms SA4D both quantitatively and qualitatively.
@@ 13 Further Scene Editing Examples
SADG (Ours) | 0.8839 | 0.9833 | 0.9734 | 0.9889 | 0.9719 | 0.9976 | 0.8941 | 0.9969 | 0.9308 | 0.9917 |
@@ 14 Visualization of the Rendered View, Rendered Segmentation, Segmented Object
We also provide the visualizations of the rendered novel view, the rendered segmentation, and the segmented objects in Fig. 18, Fig. 19, and Fig. 20. The rendered segmentation of SADG is based on Gaussian clustering, which is performed in 3D space, resulting in consistent segmentation in the dynamic scene.

### v2

@@ 9.3 Google Immersive
| *TRASE (2%) | 0.8760 | 0.9829 | 0.4068 | 0.8368 | 0.9790 | 2.3117 | 0.8738 | 0.9923 | 2.2506 | 0.9154 | 0.9942 | 9.6083 | 0.9258 | 0.9904 | 1.6118 |
| TRASE (10%) | 0.8707 | 0.9828 | 1.5782 | 0.8384 | 0.9794 | 11.8430 | 0.8792 | 0.9930 | 6.1878 | 0.9160 | 0.9939 | 71.2429 | 0.9283 | 0.9908 | 6.8211 |
| TRASE (25%) | 0.8762 | 0.9832 | 3.9425 | 0.8461 | 0.9809 | 43.2682 | 0.8886 | 0.9935 | 18.8976 | 0.9180 | 0.9943 | 221.7642 | 0.9297 | 0.9907 | 19.0767 |
@@ 10 Testing Implementation Details
After the dynamic geometry reconstruction and Gaussian features are properly trained, we proceed to our GUI to select the object of interest and evaluate the model’s performance.
@@
Optional: Adjust the threshold to refine the object’s boundaries (, , , , ).
4.
@@ 11.1 Smooth Gaussian Features
As mentioned in Sec. 3.2, we utilize smooth versions of Gaussian features, which can be defined as follows:
We provide a detailed analysis of the selection of the number of nearest neighbors for the KNN algorithm in Tab. 5. Overall, when smooth Gaussian features are enabled ( and ), the performance is better than when smooth Gaussian features are disabled (). Since there is no significant improvement between and , we eventually chose for our pipeline to avoid additional overheads with the KNN sampling.
@@ 11.2 Gaussian Feature Clustering
We also employ a subsampling strategy to enable real-time Gaussian feature clustering in scenes that typically contain more than 300k Gaussians. We perform time analysis with different sub-sampling numbers of Gaussians (2%, 10%, and 25%) on RTX3080 and Intel i7-12700. For each scene, we compute the time for clustering using Python timeit package and take the average clustering time for five runs per scene. We report the average clustering time across all scenes in each benchmark and the corresponding mIoU and mAcc for our object of interest in Tab. 6. Overall, with different sub-sampling of Gaussians (2%, 10%, and 25%), the mIoU and mAcc stay almost unchanged. However, as increasing the number of Gaussians results in a longer clustering time, we believe that 2% is the optimal sub-sampling number for clustering.
11.3 Sensitivity of Post-processing Filtering Threshold
We also evaluated the sensitivity of the post-processing filtering threshold. As illustrated in Fig. 13, there exists a wide margin of effective values. This demonstrates that selecting a threshold is not difficult, as one can easily choose a valid setting by observing the performance tendency.
@@ 11.3 Sensitivity of Post-processing Filtering Threshold
There are various clustering techniques and K-means [27, 30] is one of the most widely used. K-means partitions data into clusters, where each feature belongs to the cluster with the nearest centroid (mean of the cluster). The value must be predefined, and the algorithm iteratively adjusts the centroids to minimize the sum of squared distances between features and their corresponding cluster centers. Therefore, the performance of the algorithm heavily depends on the optimal value of .
@@
| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | ||||
| Sampling rule | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
| By Area | 0.8403 | 0.9743 | 0.8171 | 0.9632 | 0.8679 | 0.9920 |
| By Score | 0.8028 | 0.9643 | 0.7998 | 0.9722 | 0.7673 | 0.9831 |
| Grid-based | 0.8721 | 0.9822 | 0.8484 | 0.9809 | 0.8708 | 0.9922 |
| Random (Ours) | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 |
11.5 Gaussian Feature Dimension
We also analyze the impact of Gaussian feature dimensionality on segmentation performance. As shown in Tab. 8, increasing does not lead to a clear improvement in mIoU or mAcc. On the other hand, the runtime costs such as storage requirements and clustering time increase significantly with larger as shown in Tab. 9. Fig. 15 visualizes these trade-offs by plotting the no-filtered mIoU from HyperNeRF-Mask against runtime costs. Ideally, the optimal configuration should be located at the top-left corner (high mIoU, low cost). The plot confirms that our chosen dimensionality () strikes the best balance between segmentation performance and computational efficiency.
| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | |||||
| D | Filtering | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
| D=16 | ✘ | 0.8730 | 0.9822 | 0.8339 | 0.9778 | 0.8702 | 0.9921 |
| ✓ | 0.8733 | 0.9822 | 0.8563 | 0.9828 | 0.9036 | 0.9945 | |
| D=32 | ✘ | 0.8760 | 0.9829 | 0.8368 | 0.9790 | 0.8738 | 0.9923 |
| ✓ (Ours) | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | |
| D=64 | ✘ | 0.8744 | 0.9825 | 0.8347 | 0.9780 | 0.8949 | 0.9939 |
| ✓ | 0.8756 | 0.9830 | 0.8573 | 0.9832 | 0.9068 | 0.9948 | |
| D=128 | ✘ | 0.8752 | 0.9826 | 0.8320 | 0.9775 | 0.8936 | 0.9938 |
| ✓ | 0.8761 | 0.9830 | 0.8573 | 0.9830 | 0.9070 | 0.9948 | |
| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | ||||
| D | Size | Time | Size | Time | Size | Time |
| D=16 | 55 | 0.17 | 168 | 0.59 | 198 | 0.69 |
| D=32 (Ours) | 67 | 0.30 | 203 | 1.10 | 239 | 1.56 |
| D=64 | 90 | 0.43 | 272 | 4.12 | 320 | 3.46 |
| D=128 | 135 | 0.82 | 410 | 8.90 | 483 | 7.06 |
11.6 Mask Sampling
We further study the influence of SAM [22] masks sampling during the feature learning stage. We test three different sampling heuristics: spatially uniform, based on the mask area (number of pixels), and based on the mask score. For the latter, we consider the stability scores from SAM [22] model, which reflects the reliability of the object against different object’s boundary changes. We have noticed many partial segments receiving high scores, especially for objects with reflective surfaces from the NeRF-DS dataset. This negatively affects the segmentation accuracy for the model that prioritizes the masks with the highest stability scores. Spatially uniform masks are retrieved by dividing the image space into 50 cells and sampling a mask per cell. If several candidates correspond to one cell, masks are selected based on their area. This method generally works well. However, true uniform sampling highly depends on mask shape, size, and arrangement, which is hard to achieve with simple heuristics.
12 Limitations and Future Work
We observe that TRASE faces difficulties in separating small objects that significantly overlap with larger ones, as illustrated in Fig. 16. Intuitively, the scale-aware training proposed by SAGA [6] should improve the segmentation of these smaller objects; however, our experiments indicate that their pipeline yields noisy segmentation results. A potential alternative is to refine the construction of the mask-based correspondence matrix . While the current implementation encodes it with binary values (0 or 1), employing soft or probabilistic values could offer greater robustness.
As described in the main paper, our method follows a two-stage training scheme. Thus, exploring the benefits of learning dynamic reconstruction and segmentation jointly is an exciting future direction of TRASE. Moreover, due to the rapid advances in the field, novel dynamic 3DGS reconstruction backbones can be combined with our soft-mining feature learning to achieve better segmentation performance.
13 Additional Evaluations of Soft-Mining Contrastive Objective
Soft vs. Hard vs. All Mining. To validate the effectiveness of our soft-mined pixel-pair sampling strategy, we compare its average intra-cluster and inter-cluster distances against two baselines: sampling all positive–negative pairs (all) and sampling only pairs that individually satisfy the thresholds (hard), as shown in Tab. 10. Our results show that soft sample mining reduces intra-cluster distances while increasing inter-cluster distances, thereby improving feature separability.
| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | ||||
| Method | intra () | inter () | intra () | inter () | intra () | inter () |
| TRASE (all) | 0.2588 | 1.3756 | 0.3251 | 1.3382 | 0.2200 | 1.3894 |
| TRASE (hard) | 0.3517 | 1.1027 | 0.3488 | 1.1078 | 0.2942 | 1.2600 |
| TRASE (Ours) | 0.2538 | 1.3761 | 0.2437 | 1.3406 | 0.1752 | 1.3925 |
Comparison to Contrastive Methods. Existing contrastive segmentation methods CGC [12] and OpenGaussian [46] treat all negative pairs equally and focus on average feature consistency. However, this often leads to ambiguous feature assignments (Fig. 17) and segmentation artifacts (Fig. 18). In contrast, TRASE optimizes per-pixel relationships by introducing soft sample mining as discussed in our main paper.
It prioritizes challenging pixels that have any negative pixel pairs that are close in feature space and any positive pairs that are far apart, leading to more discriminative feature representations. We confirm our findings by evaluating segmentation accuracy in Tab. 1 (main). Our TRASE achieves the best segmentation performance across all tested benchmarks, validating the effectiveness of the learned segmentation field.
14 Additional Feature Space Visualizations
As shown in Fig. 17, TRASE demonstrates superior performance in separating different clusters apart compared to the existing contrastive learning approaches [46, 42]. We provide additional results in Fig. 20. The features learned by TRASE are more discriminative compared to other approaches.
15 Additional Out-of-FoV Visualizations
As shown in the main paper, TRASE with soft sample mining demonstrates superior performance in the out-of-FoV viewpoint to eliminate floating artifacts compared to TRASE with hard / all sample mining and the existing contrastive learning approaches [46, 42]. We provide additional results in Fig. 18.
@@ 15 Additional Out-of-FoV Visualizations
| CGC [12] | 0.7951 | 0.9668 | 0.9163 | 0.9659 | 0.9668 | 0.9971 | 0.8443 | 0.9950 | 0.8806 | 0.9812 |
| OpenGaussian [46] | 0.7008 | 0.9458 | 0.8590 | 0.9438 | 0.8928 | 0.9908 | 0.9034 | 0.9971 | 0.8390 | 0.9694 |
| TRASE (Ours) | 0.8839 | 0.9833 | 0.9734 | 0.9889 | 0.9719 | 0.9976 | 0.8941 | 0.9969 | 0.9308 | 0.9917 |
16 Further Scene Editing Examples
We present further scene composite examples of TRASE in Fig. 24. We also compare the quality of object removal and scene composition with SA4D [19] in Fig. 25 and Fig. 26. While SA4D introduces ghostly artifacts in place of the removed objects and carries spurious Gaussians with the segmented objects, our TRASE delivers photorealistic results of the edited scene.
17 Detailed Quantitative Results per Dataset
18 Additional Qualitative Results
We provide additional qualitative results in Fig. 21 and Fig. 22 to further showcase the effectiveness of TRASE over other methods. While SAGA [7] suffers from degenerate Gaussians, our reconstruction backbone and effective post-processing mitigate the issue. At the same time, rendered masks from other methods, such as DGD [24] and SA4D [19], are inaccurate and introduce spurious regions. Our method demonstrates the best performance among all approaches. We also showcase in Fig. 23 that TRASE achieves both time and space consistency in multiview video dataset such as Neu3D.
@@ 19 Visualization of the Rendered View, Rendered Segmentation, Segmented Object
We also provide the visualizations of the rendered novel view, the rendered segmentation, and the segmented objects in Fig. 27, Fig. 28, and Fig. 29. The rendered segmentation of TRASE is based on Gaussian clustering, which is performed in 3D space, resulting in consistent segmentation in the dynamic scene.